OpenStudy (anonymous):

if x^2+x+y=8 and x>0,then what is the value of y-8/x(x+1)?

ganeshie8 (ganeshie8):

Hint : \[x^2+x = x(x+1)\]

OpenStudy (anonymous):

i did not get it?

ganeshie8 (ganeshie8):

\[x^2+x+y=8\] \[y-8 = -(x^2+x)\] divide both sides by \(x^2+x\)

OpenStudy (kittiwitti1):

Do you still need help @arisonal ?

OpenStudy (anonymous):

thank you so much i got the answer:)

OpenStudy (kittiwitti1):

... ?

OpenStudy (anonymous):

if x/3y=4/7 and z/5y=2/21,then x/z= A) 5/18 B) 18/5 C) 7/36 D)none of these

OpenStudy (anonymous):

-1